OpenStudy (anonymous):

i need helpppp -3x+9 9

OpenStudy (unklerhaukus):

division?

OpenStudy (anonymous):

yes

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {-3(x - 3) \over -3(-3)} }\) \(\Large \color{MidnightBlue}{\Rightarrow {x - 3 \over -3} }\) \(\Large \color{MidnightBlue}{\Rightarrow {-x + 3 \over 3} }\)

Parth (parthkohli):

Got it now?

OpenStudy (unklerhaukus):

\[\frac{-3x+9}{9}\]\[=\frac{3(-x+3)}{3(3)}\]\[=\frac{-x+3}{3}\]\[=\frac{-x}{3}+\frac33\]\[=1-\frac x3\]

Parth (parthkohli):

First, I factored both the numerators and denominators, then I cancelled them out.

OpenStudy (nilankshi):

thats' rite parth i was also doing the same..but u done it first..

Parth (parthkohli):

Hahaha :D

OpenStudy (anonymous):

i kinda get it, thanks (:

Parth (parthkohli):

Haha no probs :D

OpenStudy (nilankshi):

keep it up parh..;)

Parth (parthkohli):

Thanks @nilankshi

OpenStudy (nilankshi):

yr welcome parth